Grade 12th passPhysical Chemistry

the pressure of a gas at constant volume gas thermometer is 80cm of mercury in melting ice at 1atm.when the bulb is placed in aliquid, the pressure becomes 160cm of mercury.find the temperature of liquid

kalyan
6 Years agoGrade 12th pass

1 Answer

Vikas TU
6 Years ago
 
P1 = 80 cm Hg
P2 = 160 cm Hg
melting at 1 atm 
T = 0 degree C = 273.15K 
T1/T2 = P1/P2 
T2 = 160/80 * 273.15 
T2 = 546.3 K 
Hope this helps 
Good Luck